OpenStudy (anonymous):

help please? Which equation is not equivalent to the formula d = rt? r =t/d r=d/t t=d/r d=tr

OpenStudy (anonymous):

@ivettef365 could you help me with this please?

OpenStudy (ivettef365):

ok, you have d = rt, so the one that makes no sense would be r = t/d, because, to isolate r, this is the way to do it d = r t d -- -- = --- = r t t t

OpenStudy (ivettef365):

so r = d/t not t/d

OpenStudy (anonymous):

r = t/d

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

oh okay i see now, thank you (: can you help me with another please?

OpenStudy (ivettef365):

sure